OpenStudy (anonymous):

x2 – 4x – 5 = 0 What's an easier alternative to this than the quadratic equation (completing the square)?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

what two numbers multiply to make -5 and add to make -4

OpenStudy (anonymous):

can you explain further? @Thefaceless

OpenStudy (anonymous):

okay so x2-4x-5=(x-something)(x-another thing) you need to figure out what those two things are.

OpenStudy (anonymous):

for example x2-2x+1=(x-1)(x-1)

OpenStudy (anonymous):

you can distribute to check

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

I've figured out that of the equation, the x's equal 5 and -1 respectively, but instead of writing out the full equation I'm having trouble giving a more simple method to calculate it.@Thefaceless

OpenStudy (anonymous):

well the only practical methods to find solutions to a quadratics are: Quadratic formula, factor, or completing the square